OpenStudy (anonymous):

what is x^(1/3)* x^(1/3)?

OpenStudy (anonymous):

When using exponents, x^e * x^f = x^(e+f) \[x ^{1/3}= \sqrt[3]{x}; x ^{1/3}\times x ^{1/3}= x^{1/3+1/3}=x ^{2/3}\]